OpenStudy (anonymous):

The population of a town today is 49000 people. If the population decreases with a constant rate of change and the decrease is 7% in the first year, what will the town's population be in 9 years? (Round to the nearest person)

OpenStudy (anonymous):

You need to do 7 (%) x 9 = ?

OpenStudy (anonymous):

.63

OpenStudy (anonymous):

Okay, so what's 63% of 49,000?

OpenStudy (anonymous):

1813

OpenStudy (anonymous):

3087

OpenStudy (anonymous):

Almost, the answer would be 30870

OpenStudy (anonymous):

how would i solve this question

OpenStudy (anonymous):

If the population decreases with a constant percent change and the decrease is 7% in the first year, what will the town's population be in 9 years? (Round to the nearest person)

OpenStudy (anonymous):

First you would do 7 x 9 = ? That equals 63% and then you would take 63% from 49000 and round your final answer.

OpenStudy (anonymous):

18130

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, try again, do this (0.63 * 49000) =

OpenStudy (anonymous):

30870

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Would you like to go through it one more time so you fully understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, first we would do 7 x 9 which gives us 63 but we need to convert that into a decimal so 63 / 100 = 0.63 then finally we do 0.63 * 49000 = 30870

OpenStudy (anonymous):

i dont understand why my homework keeps telling me i'm wrong

OpenStudy (anonymous):

Because you need to round your answer to the nearest person as it says in the question :P

OpenStudy (anonymous):

I have to go, hope I helped.

OpenStudy (anonymous):

A meteorologist is monitoring the atmospheric pressure at various cities whose height above sea level varies. The meteorologist determines that the atmospheric pressure decreases by 12% for each additional kilometer the city is above sea level. It has been determined that the atmospheric pressure at sea level is 103 kPa (kilopascals). Determine the atmospheric pressure in the following cities. A city located 5.2 km above sea level